### v1

# Human-Object Interaction with Vision-Language Model Guided Relative Movement Dynamics

Abstract
Human-Object Interaction (HOI) is vital for advancing simulation, animation, and robotics, enabling the generation of long-term, physically plausible motions in 3D environments. However, existing methods often fall short of achieving physics realism and supporting diverse types of interactions. To address these challenges, this paper introduces a unified Human-Object Interaction framework that provides unified control over interactions with static scenes and dynamic objects using language commands. The interactions between human and object parts can always be described as the continuous stable Relative Movement Dynamics (RMD) between human and object parts. By leveraging the world knowledge and scene perception capabilities of Vision-Language Models (VLMs), we translate language commands into RMD diagrams, which are used to guide goal-conditioned reinforcement learning for sequential interaction with objects. Our framework supports long-horizon interactions among dynamic, articulated, and static objects. To support the training and evaluation of our framework, we present a new dataset named Interplay, which includes multi-round task plans generated by VLMs, covering both static and dynamic HOI tasks. Extensive experiments demonstrate that our proposed framework can effectively handle a wide range of HOI tasks, showcasing its ability to maintain long-term, multi-round transitions. For more details, please refer to our project webpage: https://rmd-hoi.github.io/.
Figure 1: Our Relative Movement Dynamics (RMD) architecture enables dynamic object interaction and long-term multi-task completion based on VLM guidance. RMD enables the automatic construction of a unified reward function applicable to various reinforcement learning interaction tasks.
1 Introduction
Methods Dynamic Object Articulated Object Automated Reward Design Long-horizon Transition HOI Goal Definition Multitask
Kinematic NSM [30] -
SAMP [8] -
SceneDiffuser [11] -
OMOMO [14] -
Physics-Based InterPhys [9] One-to-One
PhysHOI [40] Multi-to-Multi
UniHSI [45] One-to-One
RMD (Ours) Multi-to-Multi
Table 1: Comparative Analysis of Key Features between RMD and Other Methods. The proposed framework enables a unified and automated design of goal states and reward functions. Additionally, the framework facilitates fine-grained modeling of spatio-temporal human-object interactions, ensuring natural and continuous interactions with static, dynamic, and articulated objects.
Human-Object Interaction (HOI) is pivotal for progress in fields like simulation, animation and robotics. The key research goal is to generate long-term, physically plausible human motions with corresponding object movements. Existing kinematic-based data-driven approaches [64, 24, 63, 62] often struggle with artifacts and lack of realism, as they depend on static or pre-defined object trajectories that fail to capture the nuanced, interactive dynamics between humans and moving objects.
Alternatively, some recent works employ physics engines [37, 21] to create interactive environments where agents can learn interaction skills through goal-conditioned reinforcement learning [27, 56, 40, 41, 9]. These methods take advantage of the physical realism provided by simulations to train agents in a variety of tasks, resulting in more plausible motion sequences. However, existing works primarily focus on specific task designs (e.g., sitting [3, 57] or carrying [50, 33, 4, 7]), which introduces two main issues: 1) Various interaction tasks require carefully crafted, task-specific rewards designed by human experts. Due to the complexity of dynamic human-object interactions, these reward functions are extremely challenging to design [32]; 2) These task-specific training approaches require separate policy networks and struggle with long-horizon, multi-round transitions, as the sparse goal settings often leave agents in unrecoverable states for subsequent skill transitions.
Recent efforts, such as Eureka [20], have attempted to address the first issue by leveraging large language models to automate reward design and reveal latent reward structures beyond human intuition. However, the rewards generated in this manner frequently lack interpretability and do not align well with human perceptual cues, often resulting in unnatural, non-human-like motion. UniHSI [45] proposed a “chain of contacts” framework leveraging the power of Large Language Models. However, this framework focuses on “contact” as an instantaneous spatial relationship, neglecting the temporal dynamics with objects, thus making it unsuitable for handling dynamic and complex interaction tasks. Moreover, this point-goal-reaching formulation fails to fully account for the coordination of whole-body movements, resulting in suboptimal behaviors that lack the natural fluidity and human-like motion essential for realistic interactions.
Inspired by the concept of relative motion in classical mechanics [53], we propose a novel approach that models complex HOI as stable, relative movements between human and object parts. For example, when picking up a box, different body parts maintain consistent relative distances to specific points on the object, capturing not only moment-to-moment goals (e.g., approaching the box) but also transitional behaviors (e.g., moving the box with both hands). This relational constraint encapsulates the spatial, temporal, and combinatorial elements of dynamic HOI, allowing for smoother and more realistic interactions. We introduce a bipartite diagram named the Relative Movement Dynamics (RMD) to describe the spatial-temporal relationships between human parts and object parts during a sub-sequence. Benefiting from the rapid development of Vision-Language Models (VLMs) and their powerful capabilities in world knowledge, we utilize a VLM as a high-level Planner to perceive the surrounding scene and decouple text instructions into sub-sequences, automatically formatting sub-motions into the RMD bipartite diagram. Accordingly, RMD is then constructed as the target in the goal-conditioned reinforcement learning. In summary, our contributions include:
  • $\bullet$
    We propose a general physics-based HOI synthesis framework leveraging the powerful world knowledge of VLMs, supporting long-term interactions with a diverse range of objects, including static, dynamic and articulated ones.
  • $\bullet$
    We introduce Relative Movement Dynamics (RMD), a spatial-temporal bipartite diagram designed to model fine-grained, complex dynamic interactions. RMD enables the automatic construction of a unified reward function applicable to various reinforcement learning interaction tasks.
  • $\bullet$
    We create Interplay, a novel dataset including thousands of interaction plans with both static and dynamic interaction tasks. Extensive simulation results demonstrate the effectiveness of our approach.
2 Related Work
Figure 2: An overview of the concept of Relative Movement Dynamics and VLM-guided RMD Planner. RMD describes the spatial-temporal relationships between human parts and objects during a sub-sequence. We utilize a bipartite graph to formulate RMD, where each node corresponding to a human body part is uniquely connected to a node representing an object part. The edges in this graph encode the relative movement dynamics between the parts.
2.1 Human Motion Synthesis with Static Scenes
In the field of character animation and motion synthesis, most research has focused on human motion interacting with static 3D scenes [16, 56, 51, 59]. Typically, researchers break down complex instructions into several key static poses within the scene [8, 39], and then use motion inpainting and optimization techniques to generate transitions between these key poses [38]. However, this approach often leads to mediocre and inconsistent motions due to the limited expressiveness of the intermediate sequences. Recently, diffusion-based methods like [10, 11, 34, 52, 61, 36] have achieved better results in human-scene synthesis. However, these data-driven approaches are limited to generating short-term motions due to constraints in the dataset, and the physical plausibility of the generated motions is not guaranteed. Apart from data-driven kinematic approaches, some studies have explored the problem within a reinforcement learning framework. For instance,  [60, 58] achieved long-term goal-oriented behaviors by leveraging motion primitives with specific reward designs. The UniHSI [45] framework decomposes instructions into a sequence of point-reaching control tasks under a unified formulation. However, it addresses only the spatial constraints of static scenes while overlooking the temporal dynamics of objects, rendering it inadequate for managing complex dynamic interactions.
2.2 Kinematic-based Human Motion Synthesis with Dynamic Objects
Research in human motion synthesis has increasingly focused on modeling interactions with dynamic objects [31, 12, 57, 54]. Diffusion-based frameworks [15, 14, 28, 25], guide motion generation with object trajectories but often lack realism due to predefined object paths that fail to account for physical plausibility. In contrast, InterDreamer [48] first generates human motion and subsequently uses a pre-trained world model to produce object trajectories, though this approach is limited by the simplicity of the world model, leading to inaccuracies in trajectory prediction. Some studies attempt to address these limitations by jointly modeling human-object interactions with supplementary guidance techniques, including relation intervention [42], contact prediction [5], and affordance estimation [25]. Truman [13], trained on a high-quality human-scene interaction (HSI) dataset, achieves dynamic stability through an auto-regressive diffusion model guided by action labels and waypoints. Nevertheless, kinematic models continue to face challenges such as penetration, sliding issues, and difficulties in generating long-term motion, requiring extensive annotations.
2.3 Physics-based Human Motion Synthesis with Dynamic Objects
To generate physically plausible motions, reinforcement learning methods have proven effective in training human-object interaction skills based on motion capture data [3, 22, 26, 47, 49, 43]. AMP [27] introduced an adversarial motion prior framework for realistic motion synthesis. InterPhys [9] further extended this framework by incorporating an HOI motion prior, achieving success in tasks such as sitting, lying, and carrying. Further advancements have led to successful applications in sports activities, including basketball [40, 41], tennis [55], soccer [46], skating [17], and football [19, 18]. Since these approaches often depend on heuristic designs and struggle to generalize to longer, multi-round scenarios. Our method seeks to address these limitations by introducing a Relative Movement Dynamics representation, enabling the model to capture both spatial and temporal dynamics effectively. This results in physically plausible and temporally coherent interactions, eliminating the need for manual annotations and enhancing realism in dynamic, long-term HOI tasks. The detailed comparisons of key features are listed in Tab. 1.
3 Method
In the following subsections, we first illustrate the concept of the proposed Relative Movement Dynamics (Sec. 3.1). Then, we demonstrate how we use VLM to perceive the scene context and decouple the text instruction into a plan in the form of RMD (Sec. 3.2). Finally, we elaborate on the construction of the RMD controller, which receives RMD plans as input and executes them step by step (Sec. 3.3).
3.1 Relative Movement Dynamics
Complex human-object interaction (HOI) tasks can be decomposed into several sub-sequences, where each human part maintains a consistent relative movement trend with its corresponding object part until the transition to the next sub-sequence. Therefore, we universally define the objective of the interaction plan as $D$, with the formulation as,
$\mathcal{D}=\{\mathcal{G}_{0},\mathcal{G}_{1},\ldots,\mathcal{G}_{m}\},$ (1)
where $\mathcal{G}_{i}$ denotes the $i$-th plan interaction step.$\mathcal{G}$ is composed with three parts: the target location for human root, the target location for object root and the planned Relative Movement Dynamics between human parts and object parts, with the formulation as,
$\mathcal{G}=\{\mathcal{T}_{H},\mathcal{B},\mathcal{T}_{O}\},$ (2)
$\mathcal{B}$ is a bipartite graph, where each node corresponding to a human body part is uniquely connected to a node representing an object part. The edges in this graph encode the relative movement dynamics between the parts. In this context, we assume that the human body consists of 15 parts, while the object consists of $k$ parts. In this way, we can formulate each $\mathcal{B}$ as,
=$\mathcal{B}=\{e_{ij}\mid p_{h_{i}}\in P_{H},p_{o_{j}}\in P_{O},w_{ij}\in\{0,1,2,3\}\},$ (3)
where $P_{H}$ and $P_{O}$ are the sets of human body parts and object parts, respectively. The set of human body parts is given by,
$P_{H}=\{p_{h_{1}},\dots,p_{h_{15}}\},$ (4)
The set of object parts is,
$P_{O}=\{p_{o_{1}},\dots,p_{o_{k}}\},$ (5)
Each edge $e_{ij}\in\mathcal{B}$ connects a human part $p_{h_{i}}\in P_{H}$ to an object part $p_{o_{j}}\in P_{O}$, with an edge weight $w_{ij}$ representing the state of relative movement dynamics between the two parts. The interpretation of the weights is as follows,
  • $w_{ij}=0$ indicates that $p_{h_{i}}$ and $p_{o_{j}}$ remain relatively stationary with respect to each other during this period.
  • $w_{ij}=1$ indicates that $p_{h_{i}}$ and $p_{o_{j}}$ are moving towards each other during this period.
  • $w_{ij}=2$ indicates that $p_{h_{i}}$ and $p_{o_{j}}$ are moving away from each other during this period.
  • $w_{ij}=3$ indicates that $p_{h_{i}}$ and $p_{o_{j}}$ do not maintain a stable relative motion trend during this period.
3.2 VLM-guided RMD Planner
We utilize GPT-4v [1] as our planner to infer the spatial-temporal relationships between humans and objects during interactions in specific contexts. The planner translates high-level language commands $\mathcal{L}$ into detailed plans $D$. As illustrated in Fig. 2, the RMD planner takes three main inputs: a language command $\mathcal{L}$, a top-view image representing the environmental context $\mathcal{C}$, and pre-defined prompts. Leveraging GPT-4v’s [1] vision capabilities, the RMD planner perceives the surrounding environment, identifies relevant objects, and infers their parts. To facilitate this process, we employ prompt engineering to ensure that GPT-4v [1] generates structured JSON documents, which can be easily encoded. By modifying both the instructions and the environment layout, we can generate diverse sequential interactions involving both static and dynamic objects. This approach allows us to create interaction datasets for evaluation, as detailed in the experiment section.
3.3 Policy Learning via RMD
Receiving the detailed plan output by the RMD Planner and background scenarios in the form of meshes and point clouds as input, the policy outputs specific torque for each joint to activate the agent and sequentially execute the plan. We begin by introducing the preliminaries of our method. Then, we describe the detailed formulation of the policy, including how the plan is translated into the goal state and how we construct the corresponding reward to encourage the agent to perform actions that align with the goal expectations.
Preliminary. Our physics-based animation framework is built upon goal-conditioned reinforcement learning. At each time step $t$, the agent samples an action from its policy $\pi(a_{t}\mid\mathbf{s}_{t},\mathbf{g}_{t})$ based on the current state $\mathbf{s}_{t}$ and the goal state $\mathbf{g}_{t}$. After executing the action, the environment transitions to the next state $\mathbf{s}_{t+1}$, and the agent receives a task reward $r^{G}(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1})$. To train control policies that enable characters to perform high-level tasks in a natural and life-like manner, we adopt the AMP framework [27], which introduces a discriminator to encourage the character to produce behaviors similar to those in the dataset by providing a style reward $r^{S}(\mathbf{s}_{t},\mathbf{s}_{t+1})$. Combining these two reward terms, the total reward at time step $t$ is,
$r_{t}=w^{G}r^{G}\left(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1}\right)+w^{S}r^{S}\left(\mathbf{s}_{t},\mathbf{s}_{t+1}\right).$ (6)
Note that AMP, as well as most previous methods, spends time designing the task-specific goal state and task-specific reward term. In contrast, our framework supports fully automated goal state design and switching, as well as a unified reward signal.
Goal state via RMD. At time step $t$, given the $i$-th plan interaction step $\mathcal{G}_{i}=\{\mathcal{T}_{H},\mathcal{B},\mathcal{T}_{O}\}$, we first encode each edge $e_{ij}$ from $\mathcal{B}$ by collecting the corresponding absolute position $\mathbf{p}_{h_{i}}^{p}$ and absolute velocity $\mathbf{p}_{h_{i}}^{v}$ of the human part joint $\mathbf{p}_{h_{i}}$, as well as the absolute position $\mathbf{p}_{o_{j}}^{p}$ and absolute velocity $\mathbf{p}_{o_{j}}^{v}$ of the nearest point in the object part point cloud $\mathbf{p}_{o_{j}}\in\mathbb{R}^{3}$ from the simulation environment. Next, we calculate the relative position $\tilde{\mathbf{p}}_{ij}$ and the relative velocity $\tilde{\mathbf{v}}_{ij}$ for each pair of points and transform these quantities into the agent’s coordinate system. For each edge, a one-hot encoding $w^{\prime}_{ij}$ is applied to represent $w_{ij}$. The one-hot vector $w^{\prime}_{ij}$ is then concatenated with the regularized relative position $\tilde{\mathbf{p}}_{ij}$ and relative velocity $\tilde{\mathbf{v}}_{ij}$ at time step $t$, yielding the rmd state representation $\mathbf{s}_{t}^{\text{rmd}}=\{\tilde{\mathbf{p}}_{ij},\tilde{\mathbf{v}}_{ij},w^{\prime}_{ij}\}.$ Both $\mathcal{T}_{H}$ and $\mathcal{T}_{O}$ are in the form object(spatial relationship type). In our implementation, we first compute the object’s bounding box and get $s_{x}$, $s_{y}$, $s_{z}$, which correspond to the size of the object along the $x$-, $y$-, or $z$-axis, respectively. The position is then determined by shifting from the object center to the target according to the specific spatial relationship type. For instance, armchair(front) represents a position that is $(0.7s_{x},0,0)$ in the armchair’s local coordinate system. Next, we transform the positions of $\mathcal{T}_{H}$ and $\mathcal{T}_{O}$ into the agent’s coordinate system. This transformation guides both the agent’s and the object’s trajectories, yielding the destination state $\mathbf{d}_{t}=\{\mathbf{d}_{t}^{h},\mathbf{d}_{t}^{o}\}.$ To enhance the agent’s ability to perceive its surrounding local environment and avoid obstacles, a resolution-fixed height map $\mathbf{h}_{t}$ is incorporated, which records the heights of nearby object geometries aligned with the root’s facing direction [35, 45]. Furthermore, to enhance object perception, key parameters are extracted: the eight vertices of the bounding box $\mathbf{b}_{t}^{\text{ver}}$, the rotation angle $\mathbf{\theta}_{t}$, the linear velocity $\mathbf{v}_{t}$, and the angular velocity $\mathbf{w}_{t}$. These parameters define the object’s state $\mathbf{o}_{t}=\{\mathbf{b}_{t}^{\text{ver}},\mathbf{\theta}_{t},\mathbf{v}_{t},\mathbf{w}_{t}\}.$ Finally, combine $\mathbf{s}_{t}^{\text{rmd}}$, $\mathbf{d}_{t}$, $\mathbf{h}_{t}$, and $\mathbf{o}_{t}$, we obtain the complete set of goal state.
$\mathbf{g}_{t}=\left\{\mathbf{s}_{t}^{\text{rmd}},\mathbf{d}_{t},\mathbf{h}_{t},\mathbf{o}_{t}\right\}.$ (7)
Reward via RMD. The task reward $r^{G}(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1})$ is the weighted sum of the human root target reward, object root target reward, and RMD reward. When $r^{G}$ exceeds the threshold $\theta$, we switch the $\mathcal{G}_{i}$ to the next goal $\mathcal{G}_{i+1}$.
$r^{G}\left(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1}\right)=\lambda_{RMD}\cdot r_{\text{RMD}}+\lambda_{h}\cdot r_{h}^{\text{tar}}+\lambda_{o}\cdot r_{o}^{\text{tar}}.$ (8)
The human root target reward $r_{h}^{\text{tar}}$ is designed to encourage the human root position $x_{t}^{h_{\text{root}}}$ to approach the target position $x_{t}^{h_{\text{target}}}$.
$r_{h}^{\text{tar}}=\exp\left(-\left\|x_{t}^{h_{\text{root}}}-x_{t}^{h_{\text{target}}}\right\|^{2}\right).$ (9)
The object root target reward $r_{o}^{\text{tar}}$ is designed to encourage the object root position $x_{t}^{o_{\text{root}}}$ to approach the target position $x_{t}^{o_{\text{target}}}$.
$r_{o}^{\text{tar}}=\exp\left(-\left\|x_{t}^{o_{\text{root}}}-x_{t}^{o_{\text{target}}}\right\|^{2}\right).$ (10)
The RMD reward $r_{\text{RMD}}$ is the weighted sum of $r_{RMD}^{k}$, which guides the relative movement dynamics between each human part and object part to faithfully follow the plan. Specifically, the value of $r_{RMD}^{k}$ is determined by the edge weight $w_{ij}$ from $\mathcal{B}$.
$r_{\text{RMD}}=\sum_{k=1}^{15}\lambda_{k}\cdot r_{RMD}^{k}.$ (11)
This reward function guides the relative movement dynamics between human parts and object parts to faithfully follow the plan. Specifically, the behavior of $r_{RMD}^{k}$ is determined by the weight $w_{ij}$ as follows:
  • When $w_{ij}=0$, $r_{RMD}$ encourages the relative velocity between the two particles to approach zero, or the angle between their relative velocity and relative position to be approximately $0.5\pi$.
  • When $w_{ij}=1$, the goal is for a small relative distance between the particles, minimizing the angle between their relative velocity and relative position.
  • When $w_{ij}=2$, the aim is to increase the relative distance between the particles while still minimizing the relative velocity angle.
  • Finally, when $w_{ij}=3$, it indicates an unstable relative movement trend, where the particles may either approach or move away from each other. In such cases, we directly set the reward to its maximum value.
To balance different reward weights $\lambda$ in Eq. 8 and Eq. 11, we utilize adaptive weights proposed by [45]. See Appendix C of the supplementary for additional details.
Methods Completion Rate (%) $\uparrow$ Success Rate (%) $\uparrow$ Precision (cm) $\downarrow$
Carry Push Open Sit Lie Reach Carry Push Open Sit Lie Reach Carry Push Open Sit Lie Reach
UniHSI[45] - - - 58.9 23.2 97.1 - - - 94.3 81.5 97.5 - - - 5.6 12.8 2.1
$\text{AMP}^{*}$[27] 53.2 40.4 63.2 7.4 0.9 93.2 74.9 71.0 70.3 77.4 51.9 92.2 17.3 31.7 18.1 13.2 19.1 5.5
$\text{InterPhys}^{*}$[9] 67.8 47.1 83.2 23.2 2.7 95.3 88.2 78.1 90.1 88.1 76.4 95.9 12.3 24.7 10.1 10.3 14.8 3.4
Ours 88.3 84.1 91.2 92.6 62.0 97.5 93.3 90.1 95.1 95.6 87.0 97.8 10.1 18.2 5.2 4.9 9.1 1.9
multi-one 75.1 71.3 77.9 81.0 49.2 96.2 90.5 86.4 92.1 94.2 83.1 97.2 13.6 21.0 7.3 5.5 14.3 2.0
one-multi 11.7 59.2 68.7 65.8 28.1 97.2 14.3 81.2 90.8 92.3 81.6 97.4 21.3 32.2 7.2 9.1 17.2 2.0
w.o. $\tilde{\mathbf{p}}_{ij}$ 71.7 73.8 77.0 72.9 36.8 96.8 88.9 87.8 90.3 91.8 79.8 96.8 10.9 19.7 10.1 5.6 12.8 2.9
w.o. $\tilde{\mathbf{v}}_{ij}$ 78.2 76.7 75.2 76.1 42.1 96.9 90.1 84.2 87.8 92.1 79.2 96.9 11.3 20.0 9.9 6.1 14.2 2.5
w.o. $w^{\prime}_{ij}$ 69.1 68.4 62.1 67.0 30.3 95.1 82.3 82.3 86.2 90.3 78.9 95.2 14.1 23.9 9.2 7.6 18.6 2.2
Table 2: Comparison with Baselines in Single-Task Scenario. * Indicates that these baselines have been re-implemented based on their original papers or publicly available code to ensure a consistent experimental setup. Our method achieves state-of-the-art quantitative results in both static and dynamic interaction tasks.
4 Experiment
Our experiments are organized into two main sections: the first evaluates HOI skill learning in a single-task scenario (Sec. 4.1), and the second evaluates it in a long-horizon multi-task scenario (Sec. 4.2).
Implementation Details.
To ensure our agent interacts with objects in a natural way, we collect motion clips from SAMP [9], OMOMO [14], CIRCLE [2]. We conduct experiments in parallel simulated environments within IsaacGym [21], using PyTorch for neural network implementation. In line with previous studies [9, 45], our physical humanoid model consists of 15 rigid body parts and 28 joints, all controlled by a PD controller. The simulator operates at a frequency of 60 Hz, while policy updates are made at 30 Hz. We utilize Proximal Policy Optimization (PPO) [29] to train the policy network on an NVIDIA RTX 3090 GPU.
4.1 Comparisons in a Single-Task Scenario
Settings.
We select a set of interaction tasks that are commonly used in prior works, as well as some novel tasks that have not been previously reported. These include three static HOI tasks: reaching, sitting, and lying down, and three dynamic HOI tasks: carrying, pushing, and opening. While previous works [9, 45] primarily focus on the approach and interaction with objects, they often overlook a critical aspect: after interacting with an object, the agent must return to a neutral state to facilitate subsequent interactions. For instance, in the sitting task, previous methods consider the task complete once the agent is seated. However, this disregards the need for the agent to return to a standing position in order to transition smoothly to the next task. Therefore, we redefine the task completion criteria by introducing a “leaving” step: after interacting with an object, the agent must stand up and walk to a designated position. This step ensures that the agent returns to a neutral pose, which is essential for smooth transitions between tasks. In line with prior works, we use static, movable, and articulated objects from PartNet [23], 3D-Front [6], CIRCLE [2], SAMP [8], and PartNet-Mobility [44]. For each episode, objects are initialized with a random orientation (ranging from 0 to $2\pi$), a random distance (from 4 to 10 meters), and a random scale (from 0.8 to 1.2). The finish position is then randomly sampled at a point 3 meters away from the object. To ensure a fair comparison with previous state-of-the-art methods, we modify the implementations of AMP [27], InterPhys [9], and UniHSI [45] by adjusting their goal states and corresponding reward functions.
Metrics.
In line with prior works [9, 45], we report the Success Rate to measure the percentage of trials in which the character successfully interacts with the object. A trial is considered successful if the distance between the object’s root and a specific humanoid part is less than 20 cm (for tasks like reaching, sitting, or lying down) or if the distance between the object’s root and the target location is less than 20 cm (for tasks like carrying, pushing, or opening). Additionally, we report the Completion Rate, which measures the percentage of trials in which the character successfully completes the interaction with the object and returns to a neutral standing pose (with the character’s root within 20 cm of the finish position). Furthermore, we utilize Precision to quantify the average distance between a specific human body part and its corresponding target position over the course of the entire interaction in successfully completed trials. All metrics are evaluated over 4096 trials per task.
Result Analysis.
As shown in Tab. 2, our method achieves higher or comparable performance across all these metrics compared to the baselines. Our method outperforms tasks involving the recovery process of static human-object interactions, such as getting up after sitting or lying down, as the Relative Movement Dynamics guide each body part to move away from the object. In contrast, as shown in Fig. 3, UniHSI [45] struggles to get up because it treats the human-object interaction as a sequence of independent spatial reaching tasks, neglecting the temporal dynamics that coordinate the movements of different body parts. Similarly, InterPhys [9] produces unnatural and unstable motions to complete tasks, such as abrupt kicking or thrusting forward, due to its lack of fine-grained spatial-temporal guidance for coordinating the movements of different body parts. Instead, our design ensures that the agent keeps both hands and arms relatively stationary against the back of the sofa. More qualitative visualization videos can be found in the supplementary materials.
Figure 3: Visualization for Qualitative Comparison. Other methods exhibit unnatural motion (InterPhys) or incomplete interactions (UniHSI), whereas our method demonstrates human-like motion quality in qualitative assessments. More qualitative visualization videos can be found in the supplementary materials.
Methods Completion Rate (%) $\uparrow$ Sub-step Completion Ratio (%) $\uparrow$ Sub-step Precision (cm) $\downarrow$
Static Interaction Dynamic Interaction Hybrid Static Interaction Dynamic Interaction Hybrid Static Interaction Dynamic Interaction Hybrid
UniHSI[45] 31.2 - - 61.3 - - 10.2 - -
$\text{InterPhys}^{*}$[9] 11.4 47.8 27.5 37.3 61.9 54.1 13.8 18.7 16.9
Ours 75.1 71.2 53.8 86.2 84.3 71.8 7.7 13.0 11.2
RMD-LLM 62.8 53.1 39.9 81.7 78.3 67.2 8.9 15.2 13.8
Table 3: Comparison with Baselines in Long-Horizon Multi-Task Scenarios. * Indicates the use of a vanilla combination and switching of different task policies, as the original InterPhys framework cannot handle multi-task scenarios.
Ablation.
We perform an ablation study on key components to demonstrate the impact of the proposed RMD. In the multi-one setup, we simplify the object representation by removing the division into distinct parts based on geometry and interaction types. Instead, we treat the object as a single entity. This modification leads to slight performance degradation as the model loses a finer-grained understanding of object geometry. Instead, we represent the object’s pose and kinematic information using the root pose and its kinematic data. This leads to a slight degradation in model performance, as it lacks a more fine-grained perception of object geometry. In the one-multi setup, we constrain our RMD Planner to model only the dynamics of a single human part interacting with one of the multiple object parts in each planned substep. This configuration is somewhat similar to UniHSI [45], and it results in an overall decline in performance metrics since setting a target for only a single body part disrupts the overall coordination of human movement and is likely to cause the optimization process to become stuck in a local optimum. The results demonstrate that our multi-multi objective significantly enhances performance in complex and challenging interaction scenarios. Furthermore, removing kinematic relation encoding($\tilde{p}^{ij}_{t}$, $\tilde{v}^{ij}_{t}$) or dynamic encoding($w^{\prime ij}_{t}$) harms performance since these components capture the spatial-temporal relationship between the human and the object during interaction. Overall, the proposed RMD plays an important role in ensuring the humanoid interacts completely with the object: approaching, interacting, and returning to an appropriate initial state.
4.2 Comparisons in a Multi-Task Scenario
While the single-task experiments provide valuable insights into the agent’s performance in handling individual human-object interactions, they focus on isolated tasks. However, real-world applications require the agent to manage multiple tasks either concurrently or sequentially, demanding more sophisticated coordination and planning. To evaluate the robustness of our method in such complex settings, we extend our experiments to a multi-task scenario. In this setup, the agent must handle various tasks within a single plan, which challenges the scalability and adaptability of the proposed Relative Movement Dynamics framework. This extension tests the ability of RMD to manage long-horizon, multi-task interactions. To support training and evaluation in this context, we introduce a new dataset, InterPlay, which contains thousands of interaction plans involving static, dynamic, and articulated objects.
InterPlay Dataset.
We utilize high-quality 3D object assets from existing datasets, including PartNet [23], 3D-FRONT [6], SAMP [8], CIRCLE [2], and PartNet-Mobility [44]. These datasets offer a diverse and detailed collection of 3D models that form the backbone of our experimental framework. We manually verified and further annotated the parts and corresponding point clouds of these 3D assets to ensure data consistency and high fidelity. During both training and inference, we randomly sampled at least two objects for interaction, while optionally including additional objects to provide a more comprehensive scene context. For each object, we randomly assigned its size, initial position, and rotation, thereby generating a wide range of spatial layouts that mimic real-world variability. We then rendered top-view images of the layout context and employed our RMD Planner to generate detailed, context-aware plans based on the given instructions. In total, we generated 1,210 plans, encompassing a wide variety of human-object interactions—including static, dynamic, and articulated objects—with varying temporal horizons within an indoor layout.
Settings.
We split our dataset into three categories: static-interaction, dynamic-interaction, and hybrid setups. In the static-only setup, we sequentially interact with at least two static objects as planned in the dataset. In the dynamic-only setup, we sequentially interact with at least two dynamic objects or articulated objects, also following the pre-planned dataset structure. In the hybrid setup, we select at least three objects, ensuring that at least one dynamic human-object interaction task is included in the overall plan. Since UniHSI [45] supports only sequential interactions with static objects, we adapt their released prompt to generate plans based on scenarios from our InterPlay-static-only dataset. InterPhys [9] supports both dynamic and static interaction tasks, but it is a task-specific policy. For a fair comparison, we combine multiple tasks in a vanilla setup to train a multi-task policy. We use one-hot encoding to represent different tasks and switch to the next goal based on heuristic rules.
Metrics.
Following Sec. 4.1, we report the Completion Rate, which measures the percentage of trials in which the character can sequentially complete all sub-step tasks and return to a natural standing pose, with the character’s root within 20 cm of the target position. We introduce the Sub-step Completion Ratio, which measures the ratio of completed sub-steps to the total number of planned sub-steps. This metric provides a more granular view of the agent’s ability to handle complex tasks with multiple sub-goals. Additionally, we use Sub-step Precision to measure the average distance between a specific human part and its corresponding target position, or between the object root and its corresponding target position.
Result Analysis.
Our method benefits from a unified formulation of interaction and fine-grained spatial-temporal guidance for all human body parts, enabling seamless transitions to a multi-task setup. This design not only streamlines the training process by maintaining consistent representations across various tasks but also significantly improves the model’s ability to generalize across different interaction scenarios. The approach achieves state-of-the-art performance in both individual sub-step execution and long-horizon transitions, reflecting its robustness in handling intricate sequences of actions. Without such a unified formulation, the straightforward combination of different tasks [9] often leads to interference between task-specific features and inefficient training dynamics. In contrast, our integrated method preserves critical spatial and temporal information throughout the pipeline. While UniHSI [45] excels at approaching and interacting with static objects, it struggles with actions such as getting up after interaction, which notably limits its performance in long-horizon scenarios. More qualitative visualization videos can be found in the supplementary materials.
Ablation.
While both VLMs and LLMs have extensive world knowledge, VLMs offer more precise spatial awareness owing to their seamless integration of visual and textual data. To explore this further, we modified the input data by replacing top-view images with richly descriptive paragraphs that detail the objects in the scene, outline their various parts, and specify their relative spatial positions. This modification allowed us to rigorously evaluate the performance of a LLM-based planner under conditions lacking explicit visual cues. The results clearly indicate that the spatial perception capability inherent in the vision model plays a critical role in facilitating effective planning.
5 Conclusion
In this paper, we proposed a unified framework for Human-Object Interaction (HOI) that leverages the concept of Relative Movement Dynamics (RMD) to model complex spatial-temporal interactions between humans and objects. By integrating vision-language models (VLMs) for high-level planning and goal-conditioned reinforcement learning for precise task execution, our approach enables long-term, physically plausible interactions. We also introduced InterPlay, a diverse dataset designed to support training and evaluation across a wide range of static and dynamic HOI tasks. Extensive experiments demonstrate that our framework outperforms state-of-the-art methods in both single-task and multi-task scenarios, achieving higher task completion rates and greater interaction precision.
6 Acknowledgement
This work was supported by Shanghai Local College Capacity Building Program (23010503100), National Natural Science Foundation of China (No.62406195), MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration, HPC Platform of ShanghaiTech University and Shanghai Engineering Research Center of Intelligent Vision and Imaging.
Supplementary Material
Appendix A Demo Video
To complement the qualitative results discussed in the main paper, we provide several demonstration videos that offer detailed visualizations of our method. These videos not only serve to illustrate the qualitative comparisons with competing approaches but also reinforce the effectiveness of our framework across various real-world scenarios. For a comprehensive understanding, we encourage the reader to refer to the supplementary materials accompanying this paper.
Single Task Scenario. Demo4.mp4 and Demo5.mp4 present a qualitative comparison between our method and UniHSI [45] for the sitting task. In these videos, when the agent attempts to stand up from a seated position, we observe that the method employed by UniHSI relies excessively on high-frequency jitter to twist the pelvis. This excessive reliance causes the agent to become stuck in a local minimum of the goal-reaching objective, leading to unrealistic and jerky motion that deviates from natural human behavior. In contrast, our method generates smooth, human-like motion, ensuring a natural transition from sitting to standing and more faithfully replicating realistic human behavior.
Additionally, Demo6.mp4 and Demo7.mp4 compare our approach with InterPhys [9] in the door-opening task. Although InterPhys achieves strong quantitative results, its generated motion appears unrealistic. Specifically, during the door-opening task, InterPhys uses only its body to push the door open, resulting in an unnatural movement pattern. In contrast, our method employs a coordinated use of both hands, which produces a more lifelike interaction and aligns closely with typical human behavior observed in similar scenarios.
Multi Task Scenario. Demo1.mp4, Demo2.mp4, and Demo3.mp4 further demonstrate long-horizon interactions involving static, dynamic, and articulated objects within a realistic indoor furniture setting. These videos illustrate how our method robustly handles multi-task scenarios by leveraging three key strengths of our framework:
  1. 1.
    Human-like Motion: Our approach is designed to ensure that the agent avoids undesirable states while transitioning smoothly between tasks. This results in fluid and coherent motions that are consistent with natural human biomechanics.
  2. 2.
    Unified Objective Design: By employing a consistent objective design across various interaction types, our framework effectively simplifies the task of multi-task learning. This unified approach reduces the need for extensive tuning of separate objectives for different tasks.
  3. 3.
    Vision-Language Model Capabilities: The integrated Vision-Language Model (VLM) leverages its visual understanding of the spatial layout along with rich world knowledge to generate plausible and contextually relevant action plans. This capability is crucial for adapting to complex environments and ensuring realistic task execution.
Figure 4: An overview of RMD Planner pipeline.
Appendix B Details of RMD Planner
As shown in Fig. 4, receiving the top-view image of the surrounding scene context and the text instruction, the RMD Planner outputs sequential sub-step plans in the form of structured JSON form so that the output can be processed directly using python script. To ensure the RMD Planner can work as we expect, we utilize different sections of prompt while each section supports individual functions. The detailed prompt for each section is listed below:
Figure 5: Details of prompt section 1.
Scene-context Understanding. Figure 5 provides the detailed prompt which guides the planner. Note that we simplify the spatial relationships between objects into seven types: center, forward, back, left, right, up, and down, such relative orientation is defined according to the sub-coordinate system of the reference object.
Figure 6: Details of prompt section 2.
RMD Concept Definition. In Fig. 6, we introduce the concept of Relative Movement Dynamics, represented as a bipartite graph modeling interactions between human body parts and object parts over time. The graph includes nodes for human and object parts, with edges capturing relative movement trends, such as stationary, approaching, receding, or unstable dynamics.
Figure 7: Details of prompt section 3.
Specific Plan Instance. Figure 7 illustrates an instance of the output plan, demonstrating the Relative Movement Dynamics during a scenario where a humanoid bends down to grab a box with both hands. The graph details specific relationships and movement trends between body parts and object parts, showcasing the concept in action. Besides, it also plans the target position of the humanoid root and the target position of the object root.
Figure 8: Details of prompt section 4.
Idea Outline. Fig. 8 provides an overview of our approach, highlighting the expected input and output of the interaction planner. Human-object interaction tasks are broken into sub-sequences, each characterized by consistent movement dynamics until transitioning to the next stage.
Figure 9: Details of prompt section 5.
Plan Rules. Figure 9 outlines the rules that the planner must follow, underscoring their importance in prompt engineering. These rules ensure that the planner generates accurate and coherent sub-sequences by providing clear guidelines on task interactions, sequence formatting, and criteria for determining movement dynamics. By adhering to these rules, the planner effectively breaks down complex interactions into manageable steps, enhancing the quality and reliability of the generated plans.
Figure 10: Details of prompt section 6.
Figure 11: Details of prompt section 6.
Reference Example. Figures 10, 11 and 12 sequentially presents a complete example of both input and output, clearly illustrating the concept and workflow of the planner.
Appendix C Additional Details of Reinforcement Learning
Our physics-based animation framework is built upon goal-conditioned reinforcement learning. At each time step $t$, the agent samples an action from its policy $\pi(a_{t}\mid\mathbf{s}_{t},\mathbf{g}_{t})$ based on the current state $\mathbf{s}_{t}$ and the goal state $\mathbf{g}_{t}$. After executing the action, the environment transitions to the next state $\mathbf{s}_{t+1}$, and the agent receives a task reward $r^{G}(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1})$. Further details on the definitions of state, reward, as well as the training procedure, are provided below.
State Definition. The state is divided into two components: the proprioception of the humanoid and object, and the goal condition. The proprioception component forms a 276-dimensional vector, encompassing the position, rotation, velocity, and angular velocity of the humanoid joints, as well as the bounding box of the object. While the root height is measured in the global reference frame, all other components are defined relative to the local frame of the character. The goal condition includes a 9×9 ego-centric heightmap, the target positions of the humanoid and object, kinematic relation encoding, and dynamic encoding for each part, resulting in a 222-dimensional vector.
Reward Definition. Recall the RMD reward $r_{\text{RMD}}$ at time step $t$ is the weighted sum of all the parts.
$r_{\text{RMD}}=\sum_{k=1}^{15}w_{k}\cdot r_{RMD}^{k}.$ (12)
The RMD reward for the $k$-th part $r_{RMD}^{k}$ is defined as:
$r_{RMD}^{k}=\begin{cases}r_{rv}^{\text{still}},&\text{if }v_{ij}=0\\ w_{pos}\cdot r_{rp}^{\text{close}}+w_{vel}\cdot r_{rv}^{\text{close}},&\text{if }v_{ij}=1\\ w_{pos}\cdot r_{rp}^{\text{away}}+w_{vel}\cdot r_{rv}^{\text{away}},&\text{if }v_{ij}=2\\ 1,&\text{if }v_{ij}=3\end{cases}$ (13)
The term $r_{rv}^{\text{still}}$ is defined by minimizing the dot product between the relative position and relative velocity of two particles. This term encourages either the relative velocity between the two particles to approach zero or the angle between their relative velocity and relative position to approach $\frac{\pi}{2}$.
$r_{rv}^{\text{still}}=\exp\left(-\left\lVert\tilde{\mathbf{p}}^{ij}_{t}\cdot\tilde{\mathbf{v}}^{ij}_{t}\right\rVert_{2}\right)$ (14)
The term $r_{rp}^{\text{close}}$ is defined to minimize the $\ell_{2}$ norm of the relative position between the two particles, thereby encouraging them to move closer.
$r_{rp}^{\text{close}}=\exp\left(-\left\lVert\tilde{\mathbf{p}}^{ij}_{t}\right\rVert_{2}\right)$ (15)
The term $r_{rv}^{\text{close}}$ encourages the projection of the relative velocity onto the direction of the relative position to approach a threshold. Note that the directions of these vectors are always opposite.
$r_{rv}^{\text{close}}=\exp\left(-\left\lVert\tilde{\mathbf{v}}^{ij}_{t}\cdot\frac{\tilde{\mathbf{p}}^{ij}_{t}}{\left\lVert\tilde{\mathbf{p}}^{ij}_{t}\right\rVert}+\epsilon\right\rVert_{2}\right)$ (16)
The term $r_{rp}^{\text{away}}$ is defined to maximize the $\ell_{2}$ norm of the relative position between the two particles, promoting a greater separation between them.
$r_{rp}^{\text{away}}=1-\exp\left(-\left\lVert\tilde{\mathbf{p}}^{ij}_{t}\right\rVert_{2}\right)$ (17)
The term $r_{rv}^{\text{away}}$ encourages the projection of the relative velocity onto the direction of the relative position to approach a threshold. Here, the directions of these vectors are always aligned.
$r_{rv}^{\text{away}}=\exp\left(-\left\lVert\tilde{\mathbf{v}}^{ij}_{t}\cdot\frac{\tilde{\mathbf{p}}^{ij}_{t}}{\left\lVert\tilde{\mathbf{p}}^{ij}_{t}\right\rVert}-\epsilon\right\rVert_{2}\right)$ (18)
Figure 12: Details of prompt section 6.
Figure 13: Visualization of long-term interaction with objects in an indoor home setting (part 1).
Figure 14: Visualization of long-term interaction with objects in an indoor home setting (part 2).
Figure 15: Visualization of long-term interaction with objects in an indoor home setting (part 3).
Figure 16: Visualization of long-term interaction with objects in an indoor home setting (part 4).
Figure 17: Visualization of long-term interaction with objects in an indoor home setting (part 5).
Figure 18: Visualization of long-term interaction with objects in an indoor home setting (part 6).
Figure 19: Visualization of long-term interaction with objects in an indoor home setting (part 7).
Figure 20: Visualization of long-term interaction with objects in an indoor home setting (part 8).