### v2

# Human-Object Interaction via Automatically Designed VLM-Guided Motion Policy

Abstract
Human-object interaction (HOI) synthesis is crucial for applications in animation, simulation, and robotics. However, existing approaches either rely on expensive motion capture data or require manual reward engineering, limiting their scalability and generalizability. In this work, we introduce the first unified physics-based HOI framework that leverages Vision-Language Models (VLMs) to enable long-horizon interactions with diverse object types—including static, dynamic, and articulated objects. We introduce VLM-Guided Relative Movement Dynamics (RMD), a fine-grained spatio-temporal bipartite representation that automatically constructs goal states and reward functions for reinforcement learning. By encoding structured relationships between human and object parts, RMD enables VLMs to generate semantically grounded, interaction-aware motion guidance without manual reward tuning. To support our methodology, we present Interplay, a novel dataset with thousands of long-horizon static and dynamic interaction plans. Extensive experiments demonstrate that our framework outperforms existing methods in synthesizing natural, human-like motions across both simple single-task and complex multi-task scenarios. For more details, please refer to our project webpage: https://vlm-rmd.github.io/.
1 Introduction
Human-Object Interaction (HOI) understanding is fundamental to advancing embodied AI systems in simulation, animation, and robotics. Existing approaches to HOI learning generally fall into two categories. The first paradigm employs motion tracking policies to imitate reference demonstrations [1, 2, 3]. While these methods can reproduce specific interaction patterns, they face critical limitations: Their heavy reliance on high-quality motion capture data creates scalability bottlenecks, and their strict adherence to reference trajectories inherently restricts their ability to generate novel interactions beyond the training distribution.
The second paradigm adopts a task-centric perspective, developing specialized policies for individual interactions like sitting [4, 5] or carrying [6, 7, 8, 9]. However, these methods encounter two fundamental challenges. First, they require labor-intensive reward engineering by domain experts - a particularly demanding requirement given the complex dynamics and contact-rich nature of realistic HOI scenarios [10]. This manual reward design process inherently limits generalizability across interaction types, creating a need for automated objective formulation. Second, constrained by artificially designed, single-objective reward mechanisms, trained policies often overfit to specific behavioral patterns. Producing task-compliant but biomechanically unrealistic motions that violate natural human kinematics.
Recent efforts such as Eureka [11] and Grove [12] leverage large language models (LLMs) to automatically generate reward function code. However, they rely on iterative search paradigms that are often sample-inefficient and computationally costly. UniHSI [13] advances this line of work by unifying reward design through a “chain-of-contacts” abstraction, which models an interaction as a discrete sequence of contact events. While conceptually elegant, this formulation overlooks the motion dynamics of interactions and lacks the capacity to model whole-body coordination or accommodate dynamic objects. Consequently, it struggles to handle dynamic interactions and often produces jittery, suboptimal behaviors even in static interaction scenarios.
Figure 1: Our framework automatically constructs goal states and reward functions for a variety of interaction tasks in reinforcement learning. Guided by VLMs, the resulting motion policy enables physics-based characters to perform long-horizon interactions with both static and dynamic objects.
To address these limitations, we propose a physics-based HOI framework that leverages VLMs to automatically construct goal states and reward functions, guiding the motion policy to perform long-term interactions with diverse objects. Specifically, we aim to empower VLM to provide motion-level, interaction-aware guidance by leveraging its abilities in motion imagination and semantic reasoning. Drawing inspiration from the classical notion of relative motion [14], we introduce Relative Movement Dynamics (RMD), a structured representation that encodes the fine-grained spatio-temporal relationships between human parts and object parts across an interaction sub-sequence. For example, in the task of lifting a box, the relative spatial configuration between the hands and the box remains stable, providing a natural constraint that captures both discrete interaction goals (e.g., contact) and continuous dynamics (e.g., coordinated movement). RMD is designed to prompt VLMs to imagine interaction dynamics while effectively grounding their high-level reasoning in motion-level patterns. This formulation enables VLMs to move beyond symbolic planning and participate in dynamic skill composition, thereby supporting more expressive and generalizable HOI motion policy learning. Building on RMD, our framework automatically constructs unified goal states that support both static and dynamic interactions, along with automatically designed reward functions that guide motion patterns consistent with the plan.
Our framework supports long-horizon interactions with diverse object types. However, existing datasets typically focus either on static interactions [13] or on object rearrangement [6], and thus fail to cover this setting. To address this gap, we construct a new dataset, Interplay, which includes long-horizon static and dynamic interaction tasks across varied scene contexts. This dataset enables systematic evaluation of our framework in a long-horizon, multi-task scenario. Experimental results demonstrate that our method achieves promising quantitative performance and generates natural, human-like motions across a diverse range of HOI tasks.
In summary, our contributions include:
  • We propose the first unified physics-based HOI synthesis framework leveraging the powerful world knowledge of VLMs, supporting long-term human-object interactions with a diverse range of objects, including static, dynamic and articulated ones.
  • We introduce VLM-Guided Relative Movement Dynamics (RMD), a fine-grained spatial-temporal bipartite diagram to automatically construct goal states and reward functions for reinforcement learning. This approach bypasses manual reward engineering while supporting diverse HOI, including both static and dynamic interactions.
  • We introduce Interplay, a novel dataset comprising thousands of interaction plans that cover long-horizon static and dynamic interaction tasks. Extensive simulation results demonstrate the effectiveness of our approach on both single-task and long-horizon multi-task scenarios.
2 Related Work
Table 1: Comparative analysis of key features between Ours and other methods.
Methods Articulated Object Automated Reward Design Dynamic Object Long-horizon Transition HOI Guidance Granularity Multitask High-level Planning Unified Policy
InterPhys [15] Coarse
InterScene [16] Rule-based FSM Coarse
TokenHSI [17] Rule-based FSM Coarse
UniHSI [13] Coarse
Ours Fine
Human Motion Synthesis with Static Scenes.
In the field of character animation and motion synthesis, most research has focused on human motion interacting with static 3D scenes [18, 19, 20, 21, 22, 23, 24, 25, 26]. Typically, researchers break down complex instructions into several key static poses within the scene [27, 28], and then use motion inpainting and optimization techniques to generate transitions between these key poses [29]. However, this approach often leads to mediocre and inconsistent motions due to the limited expressiveness of the intermediate sequences. Recently, diffusion-based methods like [30, 31, 32, 33, 34, 35] have achieved better results in human-scene synthesis. However, these data-driven approaches are limited to generating short-term motions due to constraints in the dataset, and the physical plausibility of the generated motions is not guaranteed. Apart from data-driven kinematic approaches, some studies have explored the problem within a reinforcement learning framework. For instance,  [36, 37] achieved long-term goal-oriented behaviors by leveraging motion primitives with specific reward designs. The UniHSI [13] framework decomposes instructions into a sequence of point-reaching control tasks under a unified formulation. However, it addresses only the spatial constraints of static scenes while overlooking the temporal dynamics of objects, rendering it inadequate for managing complex dynamic interactions.
Kinematic-based Human Motion Synthesis with Dynamic Objects.
Research in human motion synthesis has increasingly focused on modeling interactions with dynamic objects [38, 39, 5, 40]. Diffusion-based frameworks [41, 42, 43, 44], guide motion generation with object trajectories but often lack realism due to predefined object paths that fail to account for physical plausibility. In contrast, InterDreamer [45] first generates human motion and subsequently uses a pre-trained world model to produce object trajectories, though this approach is limited by the simplicity of the world model, leading to inaccuracies in trajectory prediction. Some studies attempt to address these limitations by jointly modeling human-object interactions with supplementary guidance techniques, including relation intervention [46], contact prediction [47], and affordance estimation [44]. Truman [48] and Lingo [49] trained on a high-quality human-scene interaction (HSI) dataset, achieves dynamic stability through an auto-regressive diffusion model guided by action labels or text. Nevertheless, kinematic models continue to face challenges such as penetration, sliding issues, and difficulties in generating long-term motion, requiring extensive annotations.
Physics-based Human Motion Synthesis with Dynamic Objects.
To generate physically plausible motions, reinforcement learning methods have proven effective in training human-object interaction skills based on motion capture data [4, 50, 51, 52, 3, 53]. AMP [54] introduced an adversarial motion prior framework for realistic motion synthesis. InterPhys [15] further extended this framework by incorporating an HOI motion prior, achieving success in tasks such as sitting, lying, and carrying. Further advancements have led to successful applications in sports activities, including basketball [1, 2], tennis [55], soccer [56], skating [57], and football [58, 59]. TokenHSI [17] integrates multiple skills into a single policy via a task tokenizer. These approaches often depend on heuristic designs and struggle to generalize to longer, multi-round scenarios. Our method seeks to address these limitations by introducing a Relative Movement Dynamics representation, enabling the model to capture both spatial and temporal dynamics effectively. This results in physically plausible and temporally coherent interactions, eliminating the need for manual annotations and enhancing realism in dynamic, long-term HOI tasks. The detailed comparisons of key features are listed in Tab. 1.
3 Method
Our method comprises two tightly coupled components. First, we describe how a VLM is leveraged to translate high-level instruction and environmental context into a sequence of structured interaction plans. This process is grounded in our proposed concept of RMD, which models fine-grained spatio-temporal relationships between human and object parts and serves as the foundation for interaction planning (Sec. 3.1). Then, we present a VLM-guided motion policy learning framework, where each RMD plan is automatically transformed into goal state and a corresponding reward function, enabling the agent to learn executable motion policies without manual specification (Sec. 3.2).
Figure 2: An overview of our architecture. Receiving instruction and environment context as input, the VLM-Guided RMD Planner generates a multi-step interaction plan in the form of RMD. Based on this plan, our framework automatically designs both goal states and reward functions, enabling the VLM-Guided Motion Policy to execute the interaction step by step.
3.1 VLM-Guided RMD Planner
We employ GPT-4V [60] as our VLM planner to bridge high-level task instructions and low-level interaction execution. As illustrated in Fig. 2, the planner takes three inputs: a high-level textual instruction I, a top-view contextual image of the environment C, and a set of modular prompts designed to guide the decomposition of the task into detailed interaction plans.
Effectively leveraging the motion imagination capabilities of VLMs for HOI planning necessitates a principled formulation of the interaction process. Simplistic abstractions, such as the chain-of-contacts (CoC) proposed by UniHSI [13], is inadequate: by imposing only transient point-contact constraints that are discarded immediately upon satisfaction, CoC fails to capture the evolving spatio-temporal relationships critical for modeling dynamic and coordinated interactions. In contrast, we propose a novel concept, Relative Movement Dynamics (RMD), which encodes fine-grained spatial and temporal relationships between human body parts and object components throughout the course of an interaction. This structured formulation enables the VLM to reason more effectively about the physical organization of human–object interactions and to produce structured, executable plans aligned with underlying motion dynamics. Our key insight is that human-object interactions can be abstracted as the relative motion between two sets of rigid bodies: human body parts and object parts, evolving over time. We formalize this abstraction as a bipartite graph $\mathcal{B}$ that encodes part-level motion dynamics. Let
$P_{H}=\{p_{h_{1}},p_{h_{2}},\dots,p_{h_{m}}\},\quad P_{O}=\{p_{o_{1}},p_{o_{2}},\dots,p_{o_{n}}\}$ (1)
denote the sets of human and object parts, respectively. Then the graph $\mathcal{B}$ is defined as,
$\mathcal{B}=(V,E,w),\qquad V=P_{H}\cup P_{O},\quad E\subseteq P_{H}\times P_{O},\quad w:E\to\{0,1,2,3\}.$ (2)
Each edge $e_{ij}=(p_{h_{i}},p_{o_{j}})\in E$ connects a human part $p_{h_{i}}\in P_{H}$ to an object part $p_{o_{j}}\in P_{O}$, with an associated weight $w_{ij}\in\{0,1,2,3\}$ that characterizes their relative-motion pattern: $w_{ij}=0$ indicates stationary contact, $w_{ij}=1$ denotes approaching motion, $w_{ij}=2$ represents separating motion, and $w_{ij}=3$ implies no consistent relative trend.
To facilitate the planner’s ability to infer such dynamics, we adopt a modular prompt design strategy. Each prompt segment is crafted to trigger a specific reasoning ability, such as environmental parsing, part-level object understanding, motion dynamic inference, or symbolic representation generation. Through this mechanism, the VLM is guided to produce both the RMD graph $\mathcal{B}$ and two spatial anchors: the target location for the human root $\mathcal{T}_{H}$ and the target location for the object root $\mathcal{T}_{O}$, which together define the global interaction objective.
Finally, the planner outputs an interaction plan $\mathcal{D}$ as a sequence of $N$ structured steps,
$\mathcal{D}=\{\mathcal{G}_{1},\mathcal{G}_{2},\ldots,\mathcal{G}_{N}\},$ (3)
where each step $\mathcal{G}_{i}$ is a triplet that specifies the spatial goals and dynamic pattern,
$\mathcal{G}_{i}=\{\mathcal{T}_{H},\mathcal{T}_{O},\mathcal{B}\}.$ (4)
3.2 Automatic Policy Learning via VLM-Guided RMD
Upon receiving the plan generated by the VLM-guided RMD Planner, the policy produces joint-level torques to actuate the agent and execute the plan. Our physics-based HOI policy learning is formulated within the framework of goal-conditioned reinforcement learning. At each timestep $t$, the agent samples an action from its policy $\pi(a_{t}\mid\mathbf{s}_{t},\mathbf{g}_{t})$, conditioned on the current state $\mathbf{s}_{t}$ and the goal state $\mathbf{g}_{t}$. After executing the action, the environment transitions to the next state $\mathbf{s}_{t+1}$, and the agent receives two types of rewards: a task reward $r^{G}(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1})$, which incentivizes behaviors aligned with achieving the specified goal $\mathbf{g}_{t}$, and a style reward $r^{S}(\mathbf{s}_{t},\mathbf{s}_{t+1})$, which encourages the agent to produce natural behaviors as proposed in AMP [54]. Unlike AMP and most prior methods that rely on manually specifying task-specific goal states and handcrafted reward terms, our framework supports the automatic construction of goal states and reward functions. In the following, we detail how our framework automatically generates both goal states and reward functions from VLM output, enabling the agent to learn behaviors aligned with the intended objectives.
Automatic Goal State Construction via VLM-Guided RMD.
To provide the policy with fine-grained spatio-temporal guidance during human-object interaction, we encode the $i$-th planned interaction step $\mathcal{G}_{i}=\{\mathcal{T}_{H},\mathcal{B},\mathcal{T}_{O}\}$, generated by the VLM, with both kinematic and dynamic information.
For each edge $e_{ij}\in E$, we extract from the simulator the absolute position–velocity pair of the human joint, denoted by $\mathbf{p}_{h_{i}}^{p},\mathbf{p}_{h_{i}}^{v}$, and that of the nearest surface point on the object part, $\mathbf{p}_{o_{j}}^{p},\mathbf{p}_{o_{j}}^{v}$. We then compute their relative quantities in the agent-centric coordinate frame:
$\tilde{\mathbf{p}}_{ij}=\mathbf{p}_{o_{j}}^{p}-\mathbf{p}_{h_{i}}^{p},\quad\tilde{\mathbf{v}}_{ij}=\mathbf{p}_{o_{j}}^{v}-\mathbf{p}_{h_{i}}^{v}.$ (5)
The edge weight $w_{ij}\in\{0,1,2,3\}$ is encoded as a one-hot vector $\mathbf{w}_{ij}^{\prime}\in\{0,1\}^{4}$, and concatenated with the relative position and velocity to form the feature tuple $(\tilde{\mathbf{p}}_{ij},\tilde{\mathbf{v}}_{ij},\mathbf{w}_{ij}^{\prime})$ at timestep $t$. Stacking these features across all edges yields the RMD state:
$\mathbf{s}_{t}^{\text{RMD}}=\operatorname{concat}_{(i,j)\in E}\left[\tilde{\mathbf{p}}_{ij},\;\tilde{\mathbf{v}}_{ij},\;\mathbf{w}_{ij}^{\prime}\right]\in\mathbb{R}^{\,|E|\times(3+3+4)}.$ (6)
Both $\mathcal{T}_{H}$ and $\mathcal{T}_{O}$ generated by VLM are in the form object(spatial-relationship), e.g., armchair(front). For every referenced object we approximate its geometry by an axis-aligned bounding box with edge lengths $l_{x},l_{y},l_{z}$ and centre $\mathbf{c}_{\mathrm{obj}}$. We map each spatial-relationship token $\delta$ to a displacement vector $\Delta\mathbf{q}(\delta)$ in the object’s local frame. For instance, front $\mapsto(0.7\,l_{x},0,0)$. The absolute targets for the human root and the object root are defined as,
$\mathbf{p}^{h}_{\mathrm{tar}}=\mathbf{c}_{\mathrm{obj}}+\Delta\mathbf{q}(\delta_{h}),\qquad\mathbf{p}^{o}_{\mathrm{tar}}=\mathbf{c}_{\mathrm{obj}}+\Delta\mathbf{q}(\delta_{o}).$ (7)
The destinations proposed by the VLM, originally in the global coordinate system, are transformed into the agent-centric frame to produce the final destination state,
$\mathbf{d}_{t}=\{\mathbf{d}_{t}^{h},\mathbf{d}_{t}^{o}\}.$ (8)
To improve local perception and obstacle avoidance, we attach a fixed-resolution heightmap $\mathbf{h}_{t}\!\in\!\mathbb{R}^{9\times 9}$ that records the elevation of nearby geometry in the root frame [61, 13]. In parallel, we describe object dynamics by the axis-aligned bounding-box vertices $\mathbf{V}_{t}^{\text{box}}\!\in\!\mathbb{R}^{8\times 3}$, the rotation angle $\boldsymbol{\theta}_{t}\!\in\!\mathbb{R}^{3}$, the linear velocity $\mathbf{v}_{t}\!\in\!\mathbb{R}^{3}$, and the angular velocity $\boldsymbol{\omega}_{t}\!\in\!\mathbb{R}^{3}$,
$\mathbf{o}_{t}=\bigl{\{}\mathbf{V}_{t}^{\text{box}},\;\boldsymbol{\theta}_{t},\;\mathbf{v}_{t},\;\boldsymbol{\omega}_{t}\bigr{\}}.$ (9)
By concatenating the RMD state $\mathbf{s}_{t}^{\text{RMD}}$, the destination $\mathbf{d}_{t}$, the heightmap $\mathbf{h}_{t}$, and the object state $\mathbf{o}_{t}$, we construct the complete goal representation,
$\mathbf{g}_{t}=\bigl{\{}\mathbf{s}_{t}^{\text{RMD}},\mathbf{d}_{t},\mathbf{h}_{t},\mathbf{o}_{t}\bigr{\}}.$ (10)
Automatic Reward Design via VLM-Guided RMD.
To guide human-object interaction in alignment with the $i$-th planned interaction step $\mathcal{G}_{i}=\{\mathcal{T}_{H},\mathcal{B},\mathcal{T}_{O}\}$ produced by the VLM, we design a composite reward function that promotes three key objectives: (i) guiding the human root trajectory towards the spatial target $\mathcal{T}_{H}$, (ii) guiding the object root trajectory towards $\mathcal{T}_{O}$, and (iii) enforcing the relative motion patterns prescribed by the RMD graph $\mathcal{B}$.
To realize these objectives, we define three corresponding reward terms. The term $r_{d}^{\mathrm{h}}$ encourages the human root position $x_{t}^{\mathrm{h}}$ to approach the designated target $\mathbf{d}_{t}^{h}$, and $r_{d}^{\mathrm{o}}$ encourages the object root position $x_{t}^{\mathrm{o}}$ to move toward its designated target $\mathbf{d}_{t}^{o}$,
$r_{d}^{\mathrm{h}}=\exp\left(-\left\|x_{t}^{\mathrm{h}}-\mathbf{d}_{t}^{h}\right\|^{2}\right),\qquad r_{d}^{\mathrm{o}}=\exp\left(-\left\|x_{t}^{\mathrm{o}}-\mathbf{d}_{t}^{o}\right\|^{2}\right).$ (11)
The term $r_{\text{RMD}}$ encourages each edge $e_{ij}\in\mathcal{B}$, representing a (human-part, object-part) pair, to follow the relative motion dynamics specified by its associated edge weight $w_{ij}$, as planned by the VLM. Specifically, $r_{\text{RMD}}$ is computed by summing the individual alignment reward $r_{\text{rmd}}^{ij}(\cdot)$, each of which measures how well the pair $(p_{h_{i}},p_{o_{j}})$ follows the motion pattern defined by $w_{ij}$, weighted by a coefficient $\lambda_{ij}$:
$r_{\text{RMD}}=\sum_{(i,j)\in E}\lambda_{ij}\cdot r_{\text{rmd}}\!\left(\tilde{\mathbf{p}}_{ij},\tilde{\mathbf{v}}_{ij},w_{ij}\right).$ (12)
Further implementation details of $r_{\text{rmd}}^{ij}(\cdot)$ are provided in the appendix. Combining these terms, the overall task reward at time $t$ is given by:
$r^{G}\left(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1}\right)=\lambda_{\text{RMD}}\cdot r_{\text{RMD}}+\lambda_{h}\cdot r_{d}^{\text{h}}+\lambda_{o}\cdot r_{d}^{\text{o}}.$ (13)
Once the task reward $r^{G}$, which is bounded within the range $[0,1]$, exceeds 0.9, the policy transitions to the next planned interaction step $\mathcal{G}_{i+1}$ at the following timestep. To balance the contributions of different reward terms, we adopt adaptive weighting strategies as proposed by [13], which dynamically adjust the value of $\lambda$ in Eq. 12 and Eq. 13. This design enables the current reward to be evaluated in a balanced manner and facilitates accurate assessment of progress within the current stage, thereby ensuring the appropriate timing for transitioning to the next goal.
The style reward $r^{S}\left(\mathbf{s}_{t},\mathbf{s}_{t+1}\right)$ is computed by a discriminator $D(\mathbf{s}_{t-10:t})$ based on a window of 10 consecutive states. Combining the task reward $r^{G}$ and the style reward $r^{S}$, the total reward at timestep $t$ is defined as,
$r_{t}=\alpha_{\text{task}}\,r^{G}\left(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1}\right)+\alpha_{\text{style}}\,r^{S}\left(\mathbf{s}_{t},\mathbf{s}_{t+1}\right).$ (14)
4 Experiment
We structure our experiments into two main parts: evaluating HOI skill learning in a complex long-horizon multi-task scenario (Sec. 4.1) and in a simple single-task scenario (Sec. 4.2).
InterPlay Dataset.
To evaluate our framework in a long-horizon multi-task scenario, we construct a novel dataset that includes both static and dynamic interaction tasks under diverse scene contexts. We incorporate high-quality 3D object assets from existing datasets, including PartNet [62], 3D-FRONT [63], SAMP [27], CIRCLE [64], and PartNet-Mobility [65]. Each scene contains at least two interactive objects and a minimum of two furniture assets, providing richer contextual complexity for HOI. Further details on dataset construction are provided in the appendix.
Implementation Details.
To ensure our agent interacts with objects in a natural way, we collect motion clips from SAMP [15], OMOMO [42], CIRCLE [64]. We conduct experiments in parallel simulated environments within IsaacGym [66], using PyTorch for neural network implementation. In line with previous studies [15, 13], our physical humanoid model consists of 15 rigid body parts and 28 joints, all controlled by a PD controller. The simulator operates at a frequency of 60 Hz, while policy updates are made at 30 Hz. We utilize Proximal Policy Optimization (PPO) [67] to train the policy network on an NVIDIA RTX 3090 GPU.
Table 2: Comparison with baselines in a long-horizon multi-task scenario.
Methods Completion Rate (%) $\uparrow$ Sub-step Completion Ratio (%) $\uparrow$ Sub-step Precision (cm) $\downarrow$
Static Interaction Dynamic Interaction Hybrid Static Interaction Dynamic Interaction Hybrid Static Interaction Dynamic Interaction Hybrid
$\text{InterPhys}^{*}$ [15] 21.3 47.8 27.5 37.3 61.9 54.1 13.8 18.7 16.9
$\text{TokenHSI}^{*}$ [17] 25.2 52.5 36.0 48.1 65.7 60.1 13.1 16.6 14.4
UniHSI [13] 37.2 - - 61.3 - - 10.2 - -
Ours 75.1 71.2 53.8 86.2 84.3 71.8 7.7 13.0 11.2
Ours w/ LLM 62.8 53.1 39.9 81.7 78.3 67.2 8.9 15.2 13.8
4.1 Evaluation in a Long-Horizon Multi-Task Scenario
In this setting, the agent must sequentially interact with multiple objects according to a predefined plan, posing significant challenges to the robustness of VLM-guided motion policy learning and the ability to handle long-horizon transitions.
Settings.
We split our dataset into three categories: static-interaction, dynamic-interaction, and hybrid setups. In the static-only setup, we sequentially interact with at least two static objects as planned in the dataset. In the dynamic-only setup, we sequentially interact with at least two dynamic objects or articulated objects, also following the pre-planned dataset structure. In the hybrid setup, we select at least three objects, ensuring that at least one dynamic human-object interaction task is included in the overall plan. UniHSI [13] supports only sequential interactions with static objects. We adapt its released prompt to generate plans based on scenarios from our static-only dataset. Both InterPhys [15] and TokenHSI [17] support interactions with both static and dynamic objects, but they only cover a subset of HOI tasks. To enable multi-task evaluation, we re-implement their methods by training individual skills in parallel across different environments, manually designing goal states and corresponding reward functions for each. A rule-based finite-state machine is employed to switch between tasks during execution. This adaptation requires substantial manual effort.
Metrics.
We report the Completion Rate, which measures the percentage of trials in which the character can sequentially complete all sub-step tasks and return to a natural standing pose, with the character’s root within 20 cm of the target position. We introduce the Sub-step Completion Ratio, which measures the ratio of completed sub-steps to the total number of planned sub-steps. This metric provides a more granular view of the agent’s ability to handle complex tasks with multiple sub-goals. Additionally, we use Sub-step Precision to measure the average distance between a specific human part and its corresponding target, or between the object root and its corresponding target.
Result Analysis.
As shown in Tab. 2, our approach achieves state-of-the-art performance in both individual sub-step execution and long-horizon transitions, reflecting its robustness in handling intricate sequences of actions. Our method benefits from a unified formulation of interaction and fine-grained spatial-temporal guidance for all human body parts, enabling seamless transitions to a multi-task setup. This design not only streamlines the training process by maintaining consistent representations across various tasks but also significantly improves the model’s ability to generalize across different interaction scenarios. In the absence of a unified formulation, approaches that naively combine multiple tasks, such as InterPhys [15] and TokenHSI [17], often encounter difficulties in smoothly switching between tasks. In contrast, our integrated method preserves critical spatial and temporal information throughout the pipeline. While UniHSI [13] excels at approaching and interacting with static objects, it struggles with actions such as getting up after interaction, which notably limits its performance in long-horizon scenarios. More qualitative visualization videos can be found in the supplementary materials.
Ablation.
While both VLMs and LLMs possess extensive world knowledge, VLMs exhibit superior spatial awareness and motion imagination thanks to their integration of visual and textual inputs. To evaluate this, we replaced top-view images with detailed textual descriptions and assessed an LLM-based planner under purely textual conditions. The resulting performance drop highlights the essential role of VLMs in our framework. Pretrained on diverse vision-language data, the VLM excels at generating semantically meaningful and spatially grounded plans—capabilities critical for skill learning in complex, multi-object environments. Without VLM guidance, the system struggles to produce coherent long-horizon behaviors, leading to lower task success and degraded motion quality.
Table 3: Comparison with baselines in a single-task scenario.
Methods Completion Rate (%) $\uparrow$ Success Rate (%) $\uparrow$ Precision (cm) $\downarrow$
Carry Push Open Sit Lie Reach Carry Push Open Sit Lie Reach Carry Push Open Sit Lie Reach
$\text{AMP}^{*}$ [54] 53.2 40.4 63.2 7.4 0.9 93.2 74.9 71.0 70.3 77.4 51.9 92.2 17.3 31.7 18.1 13.2 19.1 5.5
$\text{InterPhys}^{*}$ [15] 67.8 47.1 83.2 23.2 2.7 95.3 88.2 78.1 90.1 88.1 76.4 95.9 12.3 24.7 10.1 10.3 14.8 3.4
$\text{TokenHSI}^{*}$ [17] 71.2 49.3 81.1 27.8 8.9 95.7 92.2 79.5 90.8 92.8 78.2 95.9 11.2 22.5 8.7 9.8 13.2 3.5
UniHSI [13] - - - 58.9 23.2 97.1 - - - 94.3 81.5 97.5 - - - 5.6 12.8 2.1
Ours 88.3 84.1 91.2 92.6 62.0 97.5 93.3 90.1 95.1 95.6 87.0 97.8 10.1 18.2 5.2 4.9 9.1 1.9
multi-one 75.1 71.3 77.9 81.0 49.2 96.2 90.5 86.4 92.1 94.2 83.1 97.2 13.6 21.0 7.3 5.5 14.3 2.0
one-one 11.7 59.2 68.7 65.8 28.1 97.2 14.3 81.2 90.8 92.3 81.6 97.4 21.3 32.2 7.2 9.1 17.2 2.0
w.o. $\tilde{\mathbf{p}}_{ij}$ 71.7 73.8 77.0 72.9 36.8 96.8 88.9 87.8 90.3 91.8 79.8 96.8 10.9 19.7 10.1 5.6 12.8 2.9
w.o. $\tilde{\mathbf{v}}_{ij}$ 78.2 76.7 75.2 76.1 42.1 96.9 90.1 84.2 87.8 92.1 79.2 96.9 11.3 20.0 9.9 6.1 14.2 2.5
w.o. $w^{\prime}_{ij}$ 69.1 68.4 62.1 67.0 30.3 95.1 82.3 82.3 86.2 90.3 78.9 95.2 14.1 23.9 9.2 7.6 18.6 2.2
4.2 Evaluation in a Single-Task Scenario
Settings.
We select a set of interaction tasks from our InterPlay dataset, including both tasks commonly used in prior work and several novel tasks that have not been previously explored. These include three static HOI tasks: reaching, sitting, and lying down, and three dynamic HOI tasks: carrying, pushing, and opening. While previous works [15, 13] primarily focus on the approach and interaction with objects, they often overlook a critical aspect: after interacting with an object, the agent must return to a neutral state to facilitate subsequent interactions. For instance, in the sitting task, previous methods consider the task complete once the agent is seated. However, this disregards the need for the agent to return to a standing position in order to transition smoothly to the next task. Therefore, we redefine the task completion criteria by introducing a “leaving” step: after interacting with an object, the agent must stand up and walk to a designated position. This step ensures that the agent returns to a neutral pose, which is essential for smooth transitions between tasks. For each episode, objects are initialized with a random orientation (ranging from 0 to $2\pi$), a random distance (from 4 to 10 meters), and a random scale (from 0.8 to 1.2). The finish position is then randomly sampled at a point 3 meters away from the object. To ensure a fair comparison with previous state-of-the-art methods, we modify the implementations of AMP [54], InterPhys [15], TokenHSI [17], and UniHSI [13] by adjusting their goal states and corresponding reward functions.
Metrics.
We first introduce a new metric, Completion Rate, to evaluate the percentage of trials in which the character not only completes the interaction with the object but also returns to a neutral standing pose, with the root located within 20 cm of the designated finish position. This metric is designed to reflect the agent’s ability to recover from interaction and complete the full motion sequence, offering a more comprehensive evaluation of control quality. In line with prior works [15, 13], we also report the Success Rate, which measures the percentage of trials in which the character successfully interacts with the object. A trial is considered successful if the distance between the object’s root and a specific humanoid part is less than 20 cm (for tasks like reaching, sitting, or lying down), or if the distance between the object’s root and the target location is less than 20 cm (for tasks like carrying, pushing, or opening). Furthermore, we report Precision to quantify the average distance between a specific human body part and its corresponding target position over the course of the entire interaction in successfully completed trials. All metrics are evaluated over 4096 trials.
Result Analysis.
As shown in Tab. 3, our method achieves higher or comparable performance across all these metrics compared to the baselines. Our method outperforms tasks involving the recovery process of static human-object interactions, such as getting up after sitting or lying down, as the Relative Movement Dynamics guides each body part to move away from the object. In contrast, as shown in Fig. 3, UniHSI [13] struggles to get up because it treats the human-object interaction as a sequence of independent spatial reaching tasks, neglecting the temporal dynamics that coordinate the movements of different body parts. Similarly, InterPhys [15] produces unnatural and unstable motions to complete tasks, such as abrupt kicking or thrusting forward, due to its lack of fine-grained spatial-temporal guidance for coordinating the movements of different body parts. Instead, our design ensures that the agent keeps both hands and arms relatively stationary against the back of the sofa. More qualitative visualization videos can be found in the supplementary materials.
Figure 3: Visualization for qualitative comparison. Other methods exhibit unnatural motion (InterPhys) or incomplete interactions (UniHSI), whereas our method demonstrates human-like motion quality in qualitative assessments. More qualitative visualization videos can be found in the supplementary materials.
Ablation.
We perform an ablation study on key components to demonstrate the impact of the proposed VLM-guided RMD. In the multi-one setup, we simplify the object representation by removing the division into distinct parts based on geometry and interaction types. Instead, we treat the object as a single entity. This modification leads to slight performance degradation as the model loses a finer-grained understanding of object geometry. Instead, we represent the object’s pose and kinematic information using the root pose and its kinematic data. This leads to a slight degradation in model performance, as it lacks a more fine-grained perception of object geometry. In the one-one setup, we constrain our VLM-guided RMD Planner to model only the dynamics of a single human part interacting with one of the multiple object parts in each planned substep. This configuration results in an overall decline in performance metrics since setting a target for only a single body part disrupts the overall coordination of human movement and is likely to cause the optimization process to become stuck in a local optimum. The results demonstrate that our multi-multi objective significantly enhances performance in complex and challenging interaction scenarios. Furthermore, removing kinematic relation encoding($\tilde{p}^{ij}_{t}$, $\tilde{v}^{ij}_{t}$) or dynamic encoding($w^{\prime ij}_{t}$) harms performance since these components capture the spatial-temporal relationship between the human and the object during interaction. Overall, the proposed VLM-guided RMD plays an important role.
5 Conclusion
In this paper, we present a physics-based framework for synthesizing diverse Human-Object Interactions, guided by VLMs, through our proposed Relative Movement Dynamics —a spatiotemporal representation that structures part-level relationships between humans and objects. RMD automates the generation of goal states and reward functions for reinforcement learning, eliminating the need for manual engineering while supporting static, dynamic, and articulated interactions. Our method bridges high-level semantic reasoning with low-level physics-based control, producing natural motions that outperform task-specific and LLM-based baselines in both realism and adaptability. To facilitate evaluation, we introduce the Interplay dataset for long-horizon HOI tasks. Extensive results demonstrate the effectiveness of our approach on both single-task and long-horizon multi-task scenarios. Future directions include integrating diffusion models to improve motion diversity and exploring joint training of VLMs with physics-based controllers to enhance generalization and control fidelity.
6 Acknowledgement
This work was supported by National Natural Science Foundation of China (62406195, 62303319), Shanghai Local College Capacity Building Program (23010503100), ShanghaiTech AI4S Initiative SHTAI4S202404, HPC Platform of ShanghaiTech University, Core Facility Platform of Computer Science and Communication of ShanghaiTech University, and MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (ShanghaiTech University), and Shanghai Engineering Research Center of Intelligent Vision and Imaging.
Supplementary Material
[…]
Appendix
Figure A1: An overview of VLM-guided RMD Planner pipeline.
Appendix A Details of RMD Planner
As shown in Fig. A1, given a top-view image of the surrounding scene and a textual instruction, the VLM-guided RMD Planner generates a sequence of sub-step plans in the form of structured JSON. This output format facilitates direct downstream processing via Python scripts.
To ensure that the planner performs as intended, we design prompts that explicitly guide it to understand the concept of Relative Movement Dynamics (RMD), perceive the contextual scene information, decompose high-level instructions, and produce outputs in a structured format consistent with our framework. For clarity and modularity, we organize the prompt into distinct sections, each fulfilling a specific functional role as illustrated in Fig. A2 and Fig. A3. The content of each section is detailed below.
Scene-context Understanding. This section of the prompt helps the planner interpret the scene context and determine the target destination using a simplified relative spatial relationship. Specifically, we reduce spatial relations to seven canonical types: center, forward, back, left, right, up, and down.
Figure A2: Details of prompt section.
Figure A3: Details of prompt section.
These relative positions are defined with respect to a local coordinate system centered on the reference object.
RMD Concept Defination. This section of the prompt helps the planner understand the concept of RMD, clarifies its intended semantics, introduces a method for instantiating it, and provides a formal definition of its instantiated form.
Specific Plan Instance. This section presents a templated instance of an RMD-based plan, illustrating the correspondence between the template elements and RMD components, as well as how such an instance aligns with a real-world interaction scenario. This helps the planner internalize the concept and encourages outputs that conform to the desired format.
Idea Outline. This section offers a high-level overview of our approach, specifying the expected input and output of the interaction planner. Human–object interaction tasks are decomposed into a sequence of sub-steps, each characterized by consistent movement dynamics.
Plan Rules. This section enumerates the rules that the planner must follow, emphasizing their significance in prompt engineering. These rules define task interaction boundaries, output formatting, and criteria for determining movement dynamics. By adhering to these rules, the planner can effectively decompose complex interactions into manageable and coherent sub-sequences.
Reference Example. Empirical findings from the NLP community indicate that one-shot prompting often yields significantly better results than zero-shot prompting. Guided by this insight, we include a complete input–output example in the prompt to help the planner better understand the task expectations and output format.
Figure A4: Visualization of long-term interaction with objects in an indoor home setting (part 1).
Figure A5: Visualization of long-term interaction with objects in an indoor home setting (part 2).
Figure A6: Visualization of long-term interaction with objects in an indoor home setting (part 3).
Figure A7: Visualization of long-term interaction with objects in an indoor home setting (part 4).
Figure A8: Visualization of long-term interaction with objects in an indoor home setting (part 5).
Appendix B InterPlay Dataset
To evaluate our framework in long-horizon, multi-task scenarios, we construct a novel dataset, InterPlay, which encompasses both static and dynamic interaction tasks under diverse indoor scene contexts. Each data sample includes: (i) a list of processed 3D assets ready for direct use in simulation, (ii) the corresponding 3D scene layout, (iii) a top-view rendered image, and (iv) a natural language instruction describing the intended task.
In building the InterPlay dataset, we incorporate high-quality 3D object assets from several existing datasets, including PartNet [62], 3D-FRONT [63], SAMP [27], CIRCLE [64], and PartNet-Mobility [65]. From these sources, we carefully select 117 assets spanning categories such as bed, chair, sofa, washing machine, box, door, window, and wardrobe. Each asset is normalized, segmented into functional parts, and associated with part-level point clouds, based either on existing annotations or manual labeling.
To generate scene layouts, we apply a set of heuristic rules that ensure: (i) a well-distributed arrangement without object collisions, (ii) semantically meaningful placements aligned with typical household settings, and (iii) a minimum of two interactive objects and at least two furniture items per scene, thereby enriching the contextual complexity of human–object interaction. We then populate these layout templates by randomly sampling from the processed asset pool.
For each completed layout, we render a top-view image and employ GPT-4V [60] to generate a corresponding task instruction grounded in the scene context. Finally, we apply our VLM-guided RMD Planner to produce interaction plans, which are manually reviewed and refined before inclusion in the dataset.
In total, InterPlay contains 1,210 interaction plans, covering a wide range of HOI behaviors involving static, dynamic, and articulated objects with varying temporal lengths, all situated in realistic, household-inspired environments.
Appendix C Additional Details of Reinforcement Learning
Our physics-based animation framework is built upon goal-conditioned reinforcement learning. At each time step $t$, the agent samples an action from its policy $\pi(a_{t}\mid\mathbf{s}_{t},\mathbf{g}_{t})$ based on the current state $\mathbf{s}_{t}$ and the goal state $\mathbf{g}_{t}$. After executing the action, the environment transitions to the next state $\mathbf{s}_{t+1}$, and the agent receives a task reward $r^{G}(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1})$. Further details are provided below.
Action. We use a simulated humanoid character with 28 degrees of freedom (DoF), represented as a kinematic structure composed of 15 rigid bodies. The action space is defined as the target joint positions for 28 proportional–derivative (PD) controllers. At each timestep, the predicted actions specify these targets, which are then translated by the PD controllers into joint torques to drive the character’s motion.
Proprioception. The proprioceptive state $\mathbf{s}_{t}$ is a 223-dimensional feature vector that encodes the character’s physical state. It includes the positions, rotations, linear velocities, and angular velocities of all rigid bodies, expressed in the character’s local coordinate frame. The only exception is the root height, which is represented in the world coordinate frame to preserve global elevation information.
Task Reward. Recall that the term $r_{\text{RMD}}$ encourages each edge $e_{ij}\in\mathcal{B}$, representing a (human-part, object-part) pair, to follow the relative motion dynamics specified by its associated edge weight $w_{ij}$, as planned by the VLM. Specifically, $r_{\text{RMD}}$ is computed by summing the individual alignment reward $r_{\text{rmd}}^{ij}(\cdot)$, each of which measures how well the pair $(p_{h_{i}},p_{o_{j}})$ follows the motion pattern defined by $w_{ij}$, weighted by a coefficient $\lambda_{ij}$,
$r_{\text{RMD}}=\sum_{(i,j)\in E}\lambda_{ij}\cdot r_{\text{rmd}}\!\left(\tilde{\mathbf{p}}_{ij},\tilde{\mathbf{v}}_{ij},w_{ij}\right).$ (15)
Specifically, we formally define $r_{\text{rmd}}\!\left(\tilde{\mathbf{p}}_{ij},\tilde{\mathbf{v}}_{ij},w_{ij}\right)$ as,
$r_{\text{rmd}}\!\left(\tilde{\mathbf{p}}_{ij},\tilde{\mathbf{v}}_{ij},w_{ij}\right)=\begin{dcases}\exp\left(-\left(\tilde{\mathbf{p}}_{ij}\cdot\tilde{\mathbf{v}}_{ij}\right)^{2}\right),&w_{ij}=0,\\[4.0pt] \frac{1}{2}\exp\left(-\left\lVert\tilde{\mathbf{p}}_{ij}\right\rVert_{2}^{2}\right)+\frac{1}{2}\exp\left(-\left(\tilde{\mathbf{v}}_{ij}\cdot\frac{\tilde{\mathbf{p}}_{ij}}{\left\lVert\tilde{\mathbf{p}}_{ij}\right\rVert_{2}}+v^{\star}\right)^{2}\right),&w_{ij}=1,\\[4.0pt] \frac{1}{2}\left(1-\exp\left(-\left\lVert\tilde{\mathbf{p}}_{ij}\right\rVert_{2}^{2}\right)\right)+\frac{1}{2}\exp\left(-\left(\tilde{\mathbf{v}}_{ij}\cdot\frac{\tilde{\mathbf{p}}_{ij}}{\left\lVert\tilde{\mathbf{p}}_{ij}\right\rVert_{2}}-v^{\star}\right)^{2}\right),&w_{ij}=2,\\[4.0pt] 1,&w_{ij}=3.\end{dcases}$
The behavior of each reward term varies depending on the value of $w_{ij}$:
- When $w_{ij}=0$, the reward promotes either a vanishing relative velocity or an orthogonal orientation between the relative position and velocity, thereby minimizing directional alignment.
- For $w_{ij}=1$, the goal is to maintain spatial proximity while aligning the relative velocity with the direction of the relative position. In particular, the velocity projection is encouraged to approach a predefined target $v^{\star}=1\,\mathrm{m/s}$.
- When $w_{ij}=2$, the reward favors increasing separation between parts, while ensuring that the leaving speed along the direction of displacement remains close to $v^{\star}$.
- Finally, $w_{ij}=3$ denotes an unstable or unpredictable relative motion trend. Here, the reward is set to a neutral constant value of 1 to avoid imposing misleading gradients.
Training details in a single-task scenario. Our policy network is divided into a critic network and an actor network, each starting with three multilayer perceptron (MLP) layers configured with [1024, 1024, 512] units. The discriminator network follows a similar architecture, also consisting of three MLP layers with [1024, 1024, 512] units. The hyperparameters used during the training process are detailed Tab. A1,
Table A1: Hyperparameters for RMD.
Hyperparameter Value Hyperparameter Value Hyperparameter Value
Num. envs 4096 Max episode length 450 Num. epochs 25,000
Discount factor $\gamma$ 0.99 GAE parameter $\lambda$ 0.95 Minibatch size 16384
Optimizer Adam Learning rate 2e-5 Gradient clip norm 1.0
PPO horizon length 32 PPO clip $\epsilon$ 0.2 PPO miniepochs 6
Actor loss weight 1 Critic loss weight 5 Discriminator loss weight 5
AMP batch size 512 AMP sequence length $t^{*}$ 10 AMP grad. penalty $w^{gp}$ 5
Task reward weight $w^{G}$ 0.5 Disc. reward weight $w^{S}$ 0.5 Bounds loss coefficient 10
Some baseline methods either did not implement certain HOI tasks (e.g., open, push) or failed to account for post-interaction transitions with static objects (e.g., sit, lie). To enable a fair and thorough comparison, we re-implemented these baselines with carefully hand-crafted goal states and reward functions for each task. This ensures that we can comprehensively evaluate the full interaction cycle: approaching the object, interacting with it, and departing afterward, across a range of individual HOI tasks. All methods were trained using a shared set of hyperparameters to maintain consistency.
Training details in a long-horizon multi-task scenario. Our policy consists of separate actor and critic networks, each implemented as a three-layer multilayer perceptron (MLP) with hidden dimensions of [1024, 1024, 512]. The discriminator network adopts the same architecture. Both our method and the baselines are trained using a two-stage procedure:
(i) Single-task pre-training. We divide the parallel simulation environments into distinct subsets according to task type and difficulty, allowing the agent to learn basic interaction skills in a focused manner.
(ii) Long-horizon multi-task post-training. For each environment instance, tasks are randomly sampled from our InterPlay dataset, promoting the agent’s ability to handle long-horizon human–object interaction tasks in complex and diverse scenes.
To extend InterPhys to a multi-task setup, we concatenate all possible goal states and use a one-hot encoding to indicate the task identity. Additionally, we implement a conversion script to translate the sequential plans in our dataset into the format required by other baselines (i.e., sequential goal states), thereby unifying the training pipeline. During execution, a rule-based finite-state machine is used to manage task transitions in both InterPhys and TokenHSI.
Appendix D Visualizations in Supplementary Material
To complement the qualitative results presented in the main paper, we provide a demonstration video that combines the key aspects of our method. This video offers detailed visualizations showcasing the effectiveness of our framework in various real-world scenarios and compares it with competing approaches. It includes an overview of the motivation, challenges, main pipeline, and both quantitative and qualitative results. Additionally, the video highlights long-horizon interactions in both single-task and multi-task scenarios, involving static, dynamic, and articulated objects within a realistic indoor setting. In the single-task scenario, we present a comparison of our method with UniHSI [13] for the sitting task and InterPhys [15] for the door-opening task. While UniHSI struggles with high-frequency jitter and unrealistic motion, our method ensures smooth, human-like transitions. Similarly, in the door-opening task, our approach uses coordinated hand movements for a more lifelike interaction, in contrast to InterPhys, which relies solely on body movement and produces unnatural results. The multi-task scenario illustrates how our method excels by leveraging human-like motion, a unified objective design across different tasks, and the capabilities of Vision-Language Models to generate contextually relevant action plans. Additionnaly, we provide detailed visualizations alongside the Planner’s output, as shown in Fig. A4, Fig. A5, Fig. A6, Fig. A7 and Fig. A8.
Appendix E Limitations and Future Work
Limitations. While our framework advances dynamic human-object interaction modeling, it currently focuses on single-agent scenarios and does not address multi-agent interactions (e.g. collaborative tasks involving multiple humans or agents) or social dynamics (e.g. social impact considerations in shared environments). This restricts its applicability to real-world settings requiring coordination among multiple agents, such as collaborative kitchens or assistive care scenarios. Additionally, the current VLM-based planner may struggle with long-horizon tasks that demand extended temporal reasoning and hierarchical planning. For instance, tasks like cooking multi-step meals or tidying interconnected household spaces remain challenging due to the lack of explicit task decomposition mechanisms.
Future Work. To address these gaps, future research directions include extending the framework to multi-agent reinforcement learning, where agents could learn adaptive communication protocols and dynamic role allocation in socially rich environments. Simultaneously, integrating chain-of-thought prompting into the VLM-based planner could enhance its capacity for stepwise task decomposition (e.g., sequencing ingredient preparation, cooking, and serving), improving robustness in household robotics scenarios. Combining multi-agent social coordination and hierarchical reasoning would bridge the gap between atomic interactions and goal-oriented task execution, enabling scalable deployment in collaborative human environments.
Appendix F Broader Social Impact
Our method can be used to generate realistic human–object interactions for virtual avatars, simulate ergonomic behaviors, or control assistive humanoid robots in household environments. By reducing the need for hand-crafted motion design, it can accelerate the development of home robotics, virtual animation, and simulation tools. Additionally, it has the potential to improve accessibility in AR/VR applications by enabling more natural and intuitive interaction modeling, contributing to positive social impact in healthcare, rehabilitation, and inclusive technology.